# Evidence-Decision-Feedback: Theory-Driven Adaptive Scaffolding for LLM Agents

### 24 Mar 2026 v2

@@ Abstract
Multi-agent LLM architectures offer opportunities for pedagogical agents to help students construct domain knowledge and develop critical-thinking skills, yet many operate on a “one-size-fits-all” basis, limiting their ability to provide personalized support. To address this, we introduce Evidence-Decision-Feedback (EDF), a theoretical framework for adaptive scaffolding using LLMs. EDF integrates elements of intelligent tutoring systems and agentic behavior by organizing interactions around evidentiary inference, pedagogical decision-making, and adaptive feedback. We instantiate EDF through Copa, an agentic collaborative peer agent for STEM+C problem-solving. In an authentic high school classroom study, we show that EDF-guided interactions align feedback with students’ demonstrated understanding and task mastery; promote gradual scaffold fading; and support interpretable, evidence-grounded explanations without fostering overreliance.
@@ 1 Introduction
Large language model (LLM) agents afford dialogic support that aligns with learning theories emphasizing knowledge construction through social interaction, including Social Cognitive Theory [1] (SCT), social constructivist perspectives, and the Zone of Proximal Development (ZPD) [30]. However, recent studies have examined the extent to which LLMs support (or hinder) learning, with emerging findings raising concerns: Kosmyna et al. (2025) [18] reported 83% of students receiving ChatGPT assistance during essay writing were unable to quote their own work, and Zhou et al. (2025) [32] found GPT-4 feedback on compiler errors improved performance and persistence during assisted tasks but not after feedback was withdrawn. These agents can encourage overreliance, thereby impeding sustained learning and critical thinking. As such, Borchers et al. (2025) [3] concluded that LLM tutoring in its current form is “unlikely to produce learning benefits rivaling known-to-be-effective ITS tutoring.”
@@
Prior research has also shown that students often attempt to “game” intelligent learning systems by prioritizing task completion and performance metrics over deeper conceptual understanding [5]. Such behaviors can create the appearance of mastery without corresponding gains in underlying conceptual knowledge, diverging from learning-theoretic aims that emphasize sense-making, reflection, and principled scaffolding. These concerns highlight the need for pedagogical agents that do more than optimize task performance, motivating a growing call within the AIED community for LLM-based agents to support and encourage critical thinking while remaining grounded in learning theory [27, 8].
@@
Such systems must be personalized to students’ diverse needs while generating feedback interpretable to stakeholders, who may be wary of adopting systems whose reasoning is opaque or inadequately explained [11]. The AI community has recently embraced agentic AI, where agents autonomously plan actions, reason over data, and execute tasks to achieve defined goals, often within multi-agent architectures [28]. These systems are well-suited to educational contexts, as agent reasoning can support adaptation and interpretability by grounding decisions in learner-specific data and linking feedback to established learning theories. However, despite the proliferation of LLM agents in AIED research and practice, there remains a lack of LLM-era frameworks that operationalize agentic systems in systematic, pedagogically meaningful ways.
@@
There is a critical need for LLM-era frameworks that are (1) adaptive to diverse student needs, (2) effective in supporting student understanding, and (3) interpretable to stakeholders. In response, we introduce a theoretical framework for adaptive scaffolding—Evidence-Decision-Feedback (EDF; Section 3)—that draws on foundational theories including Evidence-Centered Design (ECD) [21], Stealth Assessment [26], SCT, ZPD, and social constructivist perspectives to guide the design of pedagogically grounded LLM-based agents that support student learning in real time.
@@
We developed an agentic, collaborative peer agent, Copa, to instantiate EDF by mapping each EDF component to concrete elements of Copa’s multi-agent architecture (Section 4). We deploy Copa in an authentic high school science classroom with $n=33$ dyads, where students learn physics and computing by building computational models, with the agent supporting inquiry, exploration, and critical thinking during problem-solving. Our findings suggest that EDF supports personalized, effective, and interpretable adaptive scaffolding without fostering overreliance on the agent, advancing the development of theory-aligned, agentic LLM systems in AIED.
@@ 2 Background
Before LLMs, intelligent tutoring system (ITS) frameworks like Adaptive Control of Thought–Rational (ACT-R) and Knowledge-Learning-Instruction (KLI) focused on modeling human cognition to predict task performance and align instructional design with cognitive processes. While this enabled precise model-tracing feedback, these frameworks were not well-suited to open-ended learning tasks and struggled to address the complexity and opacity of LLM systems.
Recent work has sought to improve the interpretability of LLM-based feedback by incorporating explanatory reasoning into agent responses, often using chain-of-thought (CoT), to align outputs with prompt instructions [11, 6]. While this enables LLMs to reconcile input data with instructional constraints before generating feedback, such interpretability is typically limited to surface-level explanations, leaving the underlying decision-making processes opaque.
With agentic and multi-agent architectures, decision-making can be modularized, and interactions among tools and agents can be traced back to underlying decision-making processes. Li et al.’s (2024) EduMAS framework [19] employs Concept Clarifier and Experimentation Mentor agents that collaborate to support comprehensive learning experiences. Chu et al.’s (2025) GradeOpt system [4] integrates Grader, Reflector, and Refiner agents to improve alignment with human experts. Dai et al.’s (2025) Agent4EDU framework [12] uses agentic workflows to enable targeted instructional support. However, these systems provide limited guidance on grounding agent scaffolding in pedagogy.
Many existing agents also offer limited adaptive dialogic support. Google’s LearnLM-Tutor provides scaffolding informed by stakeholder co-design [17] but does not personalize its interactions. Thomas et al. (2025) [29] ground system scaffolding in feedback theory but without dialogic interaction. Other studies similarly employ LLMs to deliver single-shot feedback, lacking the interactive and adaptive qualities central to effective dialogic learning [23, 31, 20].
The current LLM agent landscape in AIED reveals two persistent gaps: (1) a lack of principled mechanisms for personalizing scaffolding to individual learners and (2) insufficient support for interpreting how agent decisions are derived from learner evidence. While research has explored adaptive scaffolding prior to LLMs [2, 22] and in task-specific LLM applications [8], no comprehensive framework exists that addresses the following overarching question: How can we support LLM-based adaptive scaffolding that is personalized, effective, and interpretable?
@@ 3 Theoretical Framework
Figure 1: Evidence-Decision-Feedback (EDF) framework.
Over four years of participatory design (PD) with students, teachers, and learning scientists, we synthesized surveys, interviews, focus groups, pilot studies, and collaborative design activities into a framework aligned with established theory [10, 5, 7, 9, 8, 11, 13]. We present Evidence-Decision-Feedback (EDF), a principled approach to adaptive scaffolding for LLM-based pedagogical agents that addresses the limitations of existing frameworks discussed in Section 2.
EDF comprises three semi-autonomous modules that provide a blueprint for operationalizing adaptive support (shown in Figure 1). Using ECD and Stealth Assessment, the Evidence module enables stakeholders to specify the student data to be monitored to continuously update a learner model. This provides mastery evidence and is used by the Decision module to determine pedagogical intent (e.g., boosting self-efficacy) by generating a dialogue policy that is consistent with SCT and within the ZPD. The Feedback module operationalizes this policy into adaptive scaffolding, ensuring it reflects social constructivist principles. When a student responds, the response is stored as dialogue evidence and used to update the learner model before the next conversation turn.
Gray dotted arrows in Figure 1 denote interpretability, a core tenet of EDF. We define interpretability as stakeholders’ ability to understand why agent output was generated by examining how observable input data is reflected in evidentiary representations and policy explanations. Using CoT prompting, we elicit the LLM’s reasoning within each module so that agent feedback can be traced to dialogue policy selection, policy-to-evidentiary inferences, and evidence-to-input data. This provides a transparent account of how student input data is transformed into agent output, offering insight into agent decision-making.
@@ 4 Copa
To operationalize EDF, we developed Copa—a Collaborative Peer Agent for STEM+C learning. Copa is designed as a knowledgeable peer—a persona selected by students who prefer the emotional support of a peer and domain competence of an expert [5]. This role encourages inquiry, exploration, and critical thinking while minimizing the dependency risk and self-efficacy loss often associated with expert tutors. Copa is agentic, using sub-agents to gather evidence, reason about learners’ needs, and provide scaffolding within the EDF framework.
Copa reasons over student log and chat data, selects actions, and generates feedback while under constraints. This aligns with Jurenka et al.’s (2024) [17] “see what the student sees” agent design principle and prior work suggesting bounded autonomy [24]—which balances agent initiative with pedagogical guardrails, e.g., human-in-the-loop methods [13]—represents an ideal design point to support learning while preserving trust and interpretability [11, 8]. Rather than relying on unconstrained long-horizon planning, Copa’s behavior is structured around the three EDF modules. Copa’s four-agent architecture is grounded in PD, which identifies appropriate agent behavior and the data to be considered, and in prior work demonstrating that modular, multi-agent systems improve transparency and task decomposition in educational settings [19, 4, 12].
@@
Copa is embedded within the computer-based C2STEM [15] learning environment (see Figure 2). C2STEM targets 1- and 2-D kinematics: students work in dyads on various tasks, building computational models to simulate the motion of objects such as a truck (1-D) and a drone (2-D) using kinematic equations and block-based code, and interacting with Copa at their discretion. Each task is open-ended: there is no single solution, and each may be solved in myriad ways.
C2STEM interactions generate rich log data to fuel the agent’s reasoning, including logged actions (raw action sequences like adding/editing/removing blocks), model state (the code blocks on screen), and task context (the model component the student is currently addressing, e.g., variable initialization). Logged actions are mapped to processed actions to create a semantically meaningful representation for the LLM (e.g., set_block1234_vel_4 $\rightarrow$ Set Velocity to 4 m/s).
Log data also enables us to track students’ progress by defining rubric criteria that must be met for each task, even though tasks are open-ended. By translating students’ code into a canonical form, we can use the model state to determine if students have properly initialized variables, for example. Importantly, we distinguish between students’ understanding—their verbalized knowledge to the agent demonstrating what they know—and their task mastery—performance measured as the percentage score achieved for a given C2STEM task.
@@
The StrategyAgent analyzes temporal action patterns to infer cognitive problem-solving behaviors, distinguishing productive strategies (e.g., TINKERING: systematic trial-and-error) from haphazard approaches (e.g., DEPTH-FIRST ENACTING: randomly building without periodic testing). The AssessmentAgent triangulates the current model state and strategy with past learner states (e.g., EXPLORING, DEBUGGING) to generate an updated learner state. The KnowledgeAgent compares the model state to an expert reference to identify conceptual gaps within the ZPD and retrieves aligned domain knowledge (i.e., kinematics/computing information) via retrieval-augmented generation (RAG).
The DialogueAgent is the central orchestrator, integrating the Decision (green) and Feedback (purple) modules. These modules must operate synchronously, so housing them within a single agent minimizes latency, enabling asynchronous reasoning via the Evidence agents without interrupting conversational flow. Upon receiving a student query, the DialogueAgent identifies the students’ dialogue state (e.g., DEMONSTRATES_UNDERSTANDING, where students verbalize conceptual knowledge to Copa). The agent then synthesizes the dialogue state and learner model to determine pedagogical intent and generate a dialogue policy, e.g., PROBE_UNDERSTANDING (i.e., students want a direct solution or demonstrate a misconception, prompting Copa to first probe their knowledge), SUGGEST_ACTION (i.e., take a concrete action in C2STEM), or PUSH_LIMIT (i.e., push students to extend and deepen their understanding). This policy is then used by the DialogueAgent to generate a talk move, i.e., a response to students.
We evaluated LLMs from the Gemini, Claude, and GPT families and selected GPT for optimal performance. Asynchronous Evidence agents use the high-reasoning capabilities of GPT-5, while the synchronous DialogueAgent employs GPT-5-Chat to support low-latency interaction. Prompts were developed using a human-in-the-loop procedure from prior work, CoTAL [11], which combines CoT prompting with iterative refinement (active learning) to link student input data to teacher and researcher instructions. Prompts were refined across four hour-long sessions with 10-15 researchers, who interacted with Copa using C2STEM and provided feedback that informed subsequent revisions.
All agents explicitly operationalize learning theory. For example, the KnowledgeAgent estimates students’ ZPD and identifies the “immediate next step” needed based on task progress, while the DialogueAgent is instructed to “promote self-efficacy through encouragement and praise” (SCT) and ask “probing questions” (Social Constructivism). All task rubrics, as well as agent prompts, features, and their values, are detailed in our Supplementary Materials [^1] https://github.com/claytoncohn/AIED26_Supplementary_Materials.

### 29 Mar 2026 v3

@@ Abstract
LLMs offer tremendous opportunity for pedagogical agents to help students construct knowledge and develop problem-solving skills, yet many of these agents operate on a “one-size-fits-all” basis, limiting their ability to personalize support. To address this, we introduce Evidence-Decision-Feedback (EDF), a theoretical framework for adaptive scaffolding with LLM agents. EDF integrates elements of intelligent tutoring systems (ITS) and agentic behavior by organizing interactions around evidentiary inference, pedagogical decision-making, and adaptive feedback. We instantiate EDF through Copa, a Collaborative Peer Agent for STEM+C problem-solving. In an authentic high school classroom study, we show that EDF-guided interactions align feedback with students’ demonstrated understanding and task mastery; promote scaffold fading; and support interpretable, evidence-grounded explanations without fostering overreliance.
@@ 1 Introduction
Large language model (LLM) pedagogical agents afford dialogic support that aligns with learning theories emphasizing knowledge construction through social interaction, including Social Cognitive Theory [2] (SCT) and Social Constructivism [35]. However, recent studies have raised concerns: Kosmyna et al. (2025) [21] reported 83% of students receiving ChatGPT assistance while writing were unable to quote their own work, and Zhou et al. (2025) [36] found GPT-4 feedback on compiler errors benefited students during assisted tasks but not after feedback was withdrawn. While they can be helpful, these agents can also encourage overreliance [30], thereby impeding sustained learning. As such, Borchers et al. (2025) [4] concluded that LLM tutoring in its current form is “unlikely to produce learning benefits rivaling known-to-be-effective ITS tutoring.”
@@
Prior research also shows students often attempt to “game” learning systems by prioritizing task completion and performance metrics over deeper conceptual understanding [7]. Such behaviors can create the appearance of mastery without corresponding gains in conceptual knowledge, diverging from learning-theoretic aims that emphasize sense-making and self-regulation. These concerns highlight the need for pedagogical agents that do more than optimize task performance, motivating a growing call within the AIED community for LLM-based agents to support students while remaining grounded in learning theory [32, 10].
@@
Such systems must be personalized to students’ diverse needs while generating feedback interpretable to stakeholders, who may be wary of adopting systems whose reasoning is opaque [13]. The AI community has recently embraced agentic AI, where agents autonomously plan actions, reason over data, and execute tasks to achieve defined goals, often within multi-agent architectures [33]. These systems are well-suited to educational contexts, as agent reasoning can support adaptation and interpretability by grounding decisions in learner-specific data and linking feedback to learning theory. However, despite the proliferation of LLMs in the AIED community, there remains a lack of LLM-era frameworks that operationalize agentic systems in systematic, pedagogically meaningful ways.
@@
There is a critical need for LLM-era frameworks that are (1) adaptive to diverse student needs, (2) effective in supporting student understanding, and (3) interpretable to stakeholders. In response, we introduce a theoretical framework for adaptive scaffolding—Evidence-Decision-Feedback (EDF; Section 3)—that draws on foundational theories including Evidence-Centered Design (ECD) [25], Stealth Assessment [31], SCT, and social constructivist perspectives (i.e., the Zone of Proximal Development (ZPD) [35]) to guide the design of pedagogically grounded LLM-based agents that support student learning in real time.
@@
We developed a collaborative peer agent, Copa, to instantiate EDF by mapping each EDF component to concrete elements of Copa’s multi-agent architecture (Section 4). In an authentic high school science classroom with $n=33$ dyads, students learned physics and computing by building computational models in C2STEM, an open-ended learning environment, where we deployed Copa to support exploration and inquiry during problem solving [17]. Our findings suggest that EDF supports personalized, effective, and interpretable adaptive scaffolding without fostering overreliance on the agent, advancing the development of theory-aligned, agentic LLM systems in AIED.
@@ 2 Background
Table 1: Comparison of ITS and agent-based theoretical frameworks.
Framework OE LM AS SR INT LLM DC DF $T$ Primary Limitation
Classic four-model ITS architecture [27] Designed for structured tutoring settings rather than open-ended, dialogue-rich LLM use.
ACT-R / model tracing [16] Best for well-defined domains and authoring detailed cognitive steps.
KLI [20] Strong instructional guidance, less runtime tutoring decision support.
Constraint-Based Modeling [28] Good fit for constraint-based diagnosis/feedback; less so for extended, dialogic interaction.
Open Learner Models [5] Emphasizes inspectable learner representations more than end-to-end tutoring control.
Social agency frameworks [26] Emphasizes relational, conversational support more than explicit learner-model orchestration.
SRL-oriented tutoring frameworks [1] Strong for self-regulation support, but not typically framed around LLM-mediated dialogue.
LLM-based tutor architectures [23] Often prioritize flexible interaction over explicit learner modeling and theory linkage.
EDF (Ours) Connects learner modeling, learning theory, interpretability, and adaptive LLM feedback.
Legend: OE = Open-Ended; LM = Explicit Learner Model; AS = Adaptive Scaffolding; SR = Self-Regulation; INT = Interpretable/Inspectable; LLM = LLM-Ready; DC = Dialogue-Centered; DF = Dialogic Feedback (i.e., multi-turn, dynamic); TT = Theory-to-System Traceability.
Table 1 compares our proposed framework to recent alternatives across several dimensions. In contrast to intelligent tutoring systems (ITS) that emphasize direct instruction and optimized learning paths, open-ended learning environments (OELEs) emphasize exploration- and inquiry-based problem solving, in which students construct knowledge through interaction with complex scenarios (e.g., “learning-by-modeling”) rather than solely through information delivery [3].
Building effective agents for these environments necessitates robust learner modeling to support scaffolding that is both personalized to students and adaptive over time. This requires (1) attention to factors beyond cognition alone (e.g., self-regulation); (2) dialogic feedback grounded in social-cognitive and social-constructivist principles; and (3) that dialogue is treated as a first-class support mechanism, shaping what the agent says, when it probes, and how it responds. Additionally, agent decision-making and feedback must be interpretable to stakeholders and traceable through the system to its theoretical foundations.
ITS frameworks are designed primarily for well-structured domains, constrained solution paths, and tightly specified learner actions, but are less suited to OELEs where student progress is nonlinear, strategies vary, and evidence of understanding is distributed across evolving dialogue, behavior, and environment artifacts. These frameworks offer stronger theoretical grounding, interpretability, and learner modeling, but often lack dialogue-centered decision making and dialogic support for open-ended interaction. LLMs help address this gap by enabling agents to interpret richer learner evidence, engage in natural language dialogue, and generate context-sensitive feedback that adapts to students’ evolving needs. However, many existing LLM-based agents underutilize these capabilities, relying on feedback limited in personalization, adaptivity, or multi-turn interaction [19, 24, 34]. Simultaneously, advanced systems that adopt agentic or multi-agent architectures often provide less guidance on how scaffolding is grounded in pedagogy, straying from ITS theoretical foundations [22, 6, 14].
EDF addresses this trade-off by combining the interpretability and theoretical grounding of ITSs with the dialogic flexibility and dynamism of LLMs. It contributes a theory-linked orchestration framework for adaptive scaffolding in open-ended environments, in which learner-state inference, pedagogical decision making, and feedback generation are integrated to address the following question: how can we support LLM-based adaptive scaffolding that is personalized, effective, and interpretable in open-ended learning environments?
@@ 3 Theoretical Framework
Figure 1: Evidence-Decision-Feedback (EDF).
Over four years of participatory design (PD) with students, teachers, and learning scientists, we synthesized surveys, interviews, focus groups, pilot studies, and collaborative design activities into a framework aligned with established theory [8, 12, 7, 9, 11, 10, 13, 15]. We present Evidence-Decision-Feedback (EDF), a principled approach to adaptive scaffolding for LLM-based pedagogical agents that addresses the limitations of existing frameworks discussed in Section 2.
EDF comprises three semi-autonomous modules that provide a blueprint for operationalizing adaptive support (shown in Figure 1). Using ECD and Stealth Assessment, the Evidence module enables stakeholders to specify the student data to be monitored to continuously update a learner model. This provides mastery evidence and is used by the Decision module to determine pedagogical intent (e.g., boosting self-efficacy) by generating a dialogue policy that is consistent with SCT and within the ZPD. The Feedback module operationalizes this policy into adaptive scaffolding, ensuring it reflects social constructivist principles by encouraging active knowledge construction. When a student responds, the response is stored as dialogue evidence to further update the learner model.
Gray dotted arrows in Figure 1 denote interpretability, a core tenet of EDF. In our context, we define interpretability as the extent to which agent output can be traced through the input-to-evidence-to-policy-to-feedback chain using observable system artifacts, thereby supporting stakeholders’ ability to understand why the agent produced a given output. Using CoT prompting, we elicit the LLM’s reasoning within each module so that agent feedback can be traced to dialogue policy selection, policy-to-evidentiary inferences, and evidence-to-input data. This provides a transparent account of how student input data is transformed into agent output, offering insight into agent decision-making.
EDF addresses the limitations of existing frameworks by separating learner-state inference, pedagogical decision-making, and feedback generation into discrete components, linked by a shared learner representation. This makes EDF better suited to open-ended learning, where progress is nonlinear, and problems admit multiple solutions or solution paths; and to LLM-based systems, where mappings from learner evidence to agent behavior are otherwise underspecified. Rather than prescribing fixed paths or relying on opaque end-to-end generation, EDF provides a theory-linked structure for adaptive scaffolding in which each stage remains distinct, connected, and inspectable.
@@ 4 Copa
To operationalize EDF, we developed Copa—a Collaborative Peer Agent for STEM+C learning. Copa is a knowledgeable peer—a persona selected by students, who wanted Copa to offer both emotional support and domain competence [7]. This role encourages inquiry and exploration while minimizing the dependency risk and self-efficacy loss often associated with expert tutors. Copa is agentic, using sub-agents to gather evidence, reason about learners’ needs, and provide EDF-aligned scaffolding.
Copa reasons over student log and chat data, selects actions, and generates feedback while under constraints. This aligns with Jurenka et al.’s (2024) [19] “see what the student sees” design principle and prior work suggesting bounded autonomy [29]—balancing agent initiative with pedagogical guardrails—represents an ideal nexus to support learning while preserving trust [13, 10]. Copa’s behavior is structured around the three EDF modules, and its multi-agent architecture is grounded in (1) participatory design, which specifies the desired agent behavior and input data; and (2) prior work demonstrating modular, multi-agent systems can improve transparency and task decomposition in educational settings [22, 6, 14].
Copa is embedded within the C2STEM [17] learning environment (see Figure 2), which targets 1- and 2-D kinematics through open-ended dyadic computational modeling tasks. Students simulate the motion of objects such as a truck (1-D) and a drone (2-D) using kinematic equations and block-based code, interacting with Copa as they see fit. C2STEM generates rich log data to support agent reasoning, including logged actions (e.g., editing code blocks), model state (the code blocks on screen), and task context (the model component currently being addressed, e.g., variable initialization). Logged actions are further mapped to processed actions to provide the LLM with a semantically meaningful representation (e.g., set_blk1234_vel_4 $\rightarrow$ Set Velocity to 4 m/s). Log data also allows us to track students’ progress by defining rubric criteria that must be met for each task. Importantly, we distinguish between students’ understanding—their verbalized knowledge to the agent—and their task mastery—performance measured as the percentage score achieved for a given C2STEM task.
@@
The StrategyAgent analyzes temporal action patterns to infer cognitive problem-solving behaviors, distinguishing productive strategies (e.g., TINKERING: systematic trial-and-error) from haphazard approaches (e.g., DEPTH-FIRST ENACTING: randomly building without periodic testing). The AssessmentAgent triangulates the current model state and strategy with past learner states (e.g., EXPLORING, DEBUGGING) to generate an updated learner state. The KnowledgeAgent compares the model state to an expert reference to identify conceptual gaps and retrieves ZPD-aligned domain knowledge (i.e., kinematics/computing information) via retrieval-augmented generation (RAG).
The DialogueAgent is the central orchestrator that integrates the Decision (green) and Feedback (purple) modules. These modules must operate synchronously, so housing them within a single agent minimizes latency, enabling asynchronous reasoning via the Evidence agents without interrupting conversational flow. Upon receiving a student query, the DialogueAgent identifies the students’ dialogue state (e.g., DEMONSTRATES_UNDERSTANDING, where students verbalize conceptual knowledge to Copa). The agent then synthesizes the dialogue state and learner model to determine pedagogical intent via a dialogue policy, e.g., PROBE_UNDERSTANDING (i.e., students want a direct solution or demonstrate a misconception, prompting Copa to first probe their knowledge), SUGGEST_ACTION (i.e., take a concrete action in C2STEM), or PUSH_LIMIT (i.e., push students to extend and deepen their understanding). This policy is then used by the DialogueAgent to generate a talk move, i.e., a response to students.
We evaluated LLMs from the Gemini, Claude, and GPT families and selected GPT-5 for optimal performance. Prompts were developed using Chain-of-Thought Prompting + Active Learning (CoTAL) [13], a human-in-the-loop prompt engineering procedure from prior work that uses chain-of-thought (CoT) reasoning, targeted few-shot exemplars, and iterative prompt refinement to link student input data to teacher and researcher instructions. Prompts were refined across four hour-long sessions with 10-15 researchers, who interacted with Copa using C2STEM and provided feedback that informed subsequent revisions.
All agents have specific theoretical motivations. For example, the KnowledgeAgent estimates students’ ZPD and identifies the “immediate next step” needed based on task progress, while the DialogueAgent is instructed to “promote self-efficacy through encouragement and praise” (SCT) and ask “probing questions” (Social Constructivism). All task rubrics, as well as agent prompts, input features and their values, outputs, functionality, and theoretical foundations are detailed in our Supplementary Materials [^1] https://github.com/claytoncohn/AIED26_Supplementary_Materials.